# ListT5: Listwise Reranking with Fusion-in-Decoder Improves Zero-shot Retrieval

### v1

@@ 1 Introduction
Recently, listwise reranking models Ma et al. (2023); Sun et al. (2023b, a), which evaluates multiple passages together, is gaining attention for its effectiveness in zero-shot retrieval. Listwise rerankers can condition on and compare multiple passages to calibrate the relevance scores better, and can reduce the inaccuracy of predictions arising from domain shift, as theoretically supported by Xian et al. (2023), and empirically evidenced by a line of works such as Ma et al. (2023); Sun et al. (2023a).
@@
(2) Robustness to Positional Bias. In addition to efficiency, we show that ListT5 overcomes the “Lost in the middle” problem commonly encountered in LLM-based listwise rerankers Liu et al. (2023) which tend to be positionally biased to passages presented in the first and last parts of the listwise input (Sec. 4.5). We attribute the robustness of ListT5 to such bias to the nature of FiD, which distinguishes each input passage with identifiers, instead of positions as in LLMs.
(3) Zero-shot Performance. On top of efficiency and robustness, listwise reranking with ListT5 demonstrates strong performance in zero-shot retrieval, compared to not only the pointwise and pairwise methods, but also listwise methods based on LLMs that are argued to be specialized for zero-shot Pradeep et al. (2023a, b). Specifically, on the BEIR benchmark for zero-shot retrieval, our approach using T5-3B surpasses state-of-the-art methods including pointwise RankT5 (+1.4 gain in NDCG@10) (Tab. 2) and listwise RankZephyr (+1.4 gain in NDCG@10) (Tab. 3). We further perform a comprehensive ablation study and find that our key components, such as tournament sort (App. J) and generating in the increasing order of relevance, benefit zero-shot retrieval (Sec. 4.6).
@@ 2.1 Generative Models for Reranking
In the reranking scenario, rather than dual encoder models Karpukhin et al. (2020) which separately encode query and passage information, models that see query and passage information jointly at inference time Reimers and Gurevych (2019); Nogueira et al. (2020) are shown to be effective for zero-shot retrieval Rosa et al. (2022). Among those, formulating reranking as sequence generation, such as conducting listwise sorting Ma et al. (2023); Sun et al. (2023b); Pradeep et al. (2023a) or generating rationales Ferraretto et al. (2023), has shown an advantage in application to zero-shot retrieval by leveraging the language model’s auto-regressive generation capabilities. Specifically, a series of studies that use the encoder-decoder architecture of T5 (Sec. 2.2), and applying zero-shot reranking with LLMs (Sec. 2.3) has been successful.
@@ 3.1 Basic Operating Unit ($\boldsymbol{m\to r}$)
The basic operating unit of ListT5 processes a fixed number of $m$ input passages and outputs top-$r$ passages ($m\to r$), as in Fig. 2. We develop the basic unit upon the Fusion-in-Decoder (FiD) architecture Izacard and Grave (2021), which modifies the Encoder-Decoder structure of T5 Raffel et al. (2020) to process listwise inputs efficiently. Specifically, given a query $\mathbf{q}$ and a list of $m$ candidate passages $[\mathbf{p}_{1},...,\mathbf{p}_{m}]$, our basic unit concatenates each passage $\mathbf{p}_{i}$ with its index identifier $i$ and the query $\mathbf{q}$ in the following format, and then feeds them to Encoder to obtain their representation $\mathbf{h}_{i}$:
@@ 3.2 Extension with Tournament Sort ($\boldsymbol{n\to k}$)
Table 1: Comparison of the computational complexity of different ranking methods on obtaining top-$k$ passages from $n$ candidate passages. The base of the log function is $m$. While being able to see multiple passages, ListT5 has lower time complexity than pairwise models, and is competitive with pointwise models.
@@ 4.1 Overview
In this section, after explaining the training and evaluation details, we present three main results in the following order: (1) General performance and efficiency comparison of our models with respect to pointwise and listwise ranking models at Sec. 4.4, (2) Analyzing the lost-in-the-middle problem by measuring the robustness to positional bias between ListT5 and listwise ranking models at Sec. 4.5, and (3) Ablation experiments on the design choice at Sec. 4.6. All results are based on the T5-base model except explicitly mentioning 3B.
@@ 4.2 Training
To train ListT5, we use the official train subset of the MS MARCO Bajaj et al. (2018) passage ranking dataset [^1] https://microsoft.github.io/msmarco/Datasets. The dataset consists of 532,761 distinct queries and 8.8M passages, with binary annotations of relevancy. Since the ordering of relevancy between negative passages is not provided, we use a dual-encoder retrieval model, specifically COCO-DR large Yu et al. (2022), to retrieve Top-1000 passages and label relevance scores of negatives for the training dataset of MS MARCO. Experiments using GTR Ni et al. (2022), the same model used to build the training data for Zhuang et al. (2022), are reported at Appendix. B.2. For each query-positive passage pair, we randomly sample 20 groups of $m-1$ distinct negative passages. We randomly shuffle the positive and negative passages and assign identifiers $\{1,...,m\}$ to them to form each training data.
@@ 4.3 Evaluation
We measure the evaluation performance on the official dev set of MS MARCO, TREC-DL19, and TREC-DL20 Craswell et al. (2021) for in-domain performance, and the BEIR benchmark Thakur et al. (2021) for zero-shot out-domain performance. BEIR contains 18 diverse sets of domains and tasks. We download the dumped index of the Top-100 and Top-1000 retrieved passages by BM25 from the official Pyserini repository. Reranking results using advanced first-stage retrieval models (e.g., COCO-DR) are reported at App. 8. We use the evaluation metric of the Normalized Discounted Cumulative Gain@10 (NDCG@10) Järvelin and Kekäläinen (2002), the official evaluation metric for BEIR Thakur et al. (2021). All reported scores are rounded to the nearest tenth. We also report the Mean Reciprocal Rank@10 (MRR@10) scores at App. Tab. 9. For reproducibility, we include the evaluation details including details to run baseline models and links for dataset download at App. G.
@@
Efficiency in FLOPs.
@@
In addition to comparing the computational complexity in Sec. 3.2, we also conduct real-time measurements of floating point operations (FLOPs) using the FlopsProfiler of the DeepSpeed library [^2] https://deepspeed.readthedocs.io/en/latest/flops-profiler.html. We measure the FLOPs needed to rerank Top-10 passages from BM25-Top100 candidate passages for TREC-DL19 (43 queries, maximum input length of 256) on T5-base, run the profiler, and report the relative FLOPs when we set the FLOPs of MonoT5 (pointwise model) as 1333The exact value of FLOPs for MonoT5 in this setup was 229,732,539,806,400.. We also report the performance and FLOPs of the sliding window variants of ListT5, with details at App. J. The results in Fig. 4 shows that ListT5 ($r$=1) and ($r$=2) are comparable with MonoT5 and RankT5 (pointwise ranking models), and much more efficient than DuoT5 (pairwise ranking models). Also, tournament sort uses output caching to lower down FLOPs than the sliding window variants, with better performance, on the real-time FLOPs calculation (App. J). Additionally, we can parallize computation of the same level (e.g., leaf node) within query for tournament sort, while the sliding window approach should be computed sequentially. Even better, the ListT5 model is flexible - according to their needs, users can have the choice to select between the trade-off of inference speed ($r$=1) and performance ($r$=2), an option not available for other models.
4.5 Robustness to Positional Bias
Background.
Baseline Models.
Experiment and Data Setup.
4.6 Ablation Study
@@ 5 Conclusion
We proposed ListT5, a FiD-based listwise reranking model that jointly considers multiple passages as input and generates an ordered list of passages, in increasing order of relevance. ListT5 outperforms others in BEIR, with a remarkable $+$1.3 gain of average NDCG@10 performance. ListT5 is efficient, leveraging tournament sort to its advantage. Finally, we show that ListT5 overcomes the lost-in-the-middle problem in LLMs.
@@ 6 Limitations
  • We believe that the efficiency of our models can be greatly optimized using simple ideas. For example, if we can implement early stopping at sequential decoding (assuming that the model outputs permutation of 5), we can reduce the decoding step by 80% for ListT5 ($r$=1) and 60% For ListT5 ($r$=2). There could also be other options, such as designing a more efficient way that uses the tournament sort with fewer number of forward passes.
@@ B.1 Impact of the number of passages $m$ ListT5 sees at once
MS MARCO 40.7 40.7 40.5 40.7
+ Top-1000 44.7 44.9 44.6 45.0
TREC-DL19 71.2 71.8 70.1 70.5
TREC-DL20 67.3 68.1 66.9 67.2
TREC-COVID 76.7 78.3 76.2 77.9
NFCorpus 35.5 35.6 36.2 36.6
BioASQ 57.2 56.4 55.4 56.4
NQ 52.0 53.1 51.5 52.5
HotpotQA 72.1 72.6 71.4 71.9
FiQA-2018 39.5 39.6 39.0 38.9
Signal-1M (RT) 33.3 33.5 31.7 32.0
TREC-NEWS 47.9 48.5 47.1 47.8
Robust04 52.0 52.1 52.2 53.1
Arguana 49.7 48.9 38.6 46.6
Touche-2020 34.2 33.4 32.4 32.7
CQADupStack 38.4 38.8 38.2 28.8
Quora 86.1 86.4 85.5 86.8
DBPedia 43.9 43.7 42.7 43.6
SCIDOCS 17.2 17.6 17.2 18.0
FEVER 77.8 79.8 76.7 79.1
Climate-FEVER 22.8 24.0 22.7 23.9
SciFact 74.1 74.1 73.4 74.2
Avg(In-domain) 56.0 56.4 55.5 55.9
Avg(BeIR) 50.6 50.9 49.3 50.0
@@
This aligns with Sec. 4.6 from the main paper. In addition to varying the output format, we try to increase the number of passages $m$ the model sees at once, with variants on $r$ as well (See Sec. 3.1 for basic notation). The results from Tab. 6 show that setting $m$ as 10 is not as effective as setting $m$ as 5. We conjecture that there is a trade-off between the (1) benefits from seeing more negatives at once, and (2) the complexity of the ordering task itself. If we increase $m$, the model can get more information, but at the same time, the task becomes very hard, and the boundary of negative order becomes very weak (and not very informative). Therefore, in the case of T5-base models, we find that 5 performed better than 10 passages, but we believe the optimal number can be improved (or can be different) for models with different architectures or with larger parametric size.
@@ B.2 Impact of the negative selection model
Table 7: Comparison of ListT5 models trained with GTR and COCO-DR. For ListT5, the reported scores are evaluated using the (r=2) variant.

### v3

@@ 1 Introduction
Recently, listwise reranking models Ma et al. (2023); Sun et al. (2023b, a), which evaluates multiple passages together, is gaining attention for its effectiveness in zero-shot retrieval. Listwise rerankers can condition on and compare multiple passages to calibrate the relevance scores better, and can reduce the inaccuracy of zero-shot predictions arising from domain shift, as theoretically supported by Xian et al. (2023), and empirically evidenced by a line of works such as Ma et al. (2023); Sun et al. (2023a).
@@
(2) Robustness to Positional Bias. In addition to efficiency, we show that ListT5 overcomes the “Lost in the middle” problem commonly encountered in LLM-based listwise rerankers Liu et al. (2023) which tend to be positionally biased to passages presented in the first and last parts of the listwise input (Sec. 4.5) and are more robust to the change in initial ordering of the passage (Sec. 4.6). We attribute the robustness of ListT5 to such bias to the nature of FiD, which distinguishes each input passage with identifiers, instead of positions as in LLMs.
(3) Zero-shot Performance. On top of efficiency and robustness, listwise reranking with ListT5 demonstrates strong performance in zero-shot retrieval, compared to not only the pointwise and pairwise methods, but also listwise methods based on LLMs that are argued to be specialized for zero-shot Pradeep et al. (2023a, b). Specifically, on the BEIR benchmark for zero-shot retrieval, our approach using T5-3B surpasses state-of-the-art methods including pointwise RankT5 (+1.4 gain in NDCG@10) (Tab. 2) and listwise RankZephyr (+1.4 gain in NDCG@10) (Tab. 3). We further perform a comprehensive ablation study and find that our key components, such as tournament sort (App. 8) and generating in the increasing order of relevance, benefit zero-shot retrieval (Sec. 6).
@@ 2.1 Generative Models for Reranking
In the reranking scenario, rather than dual encoder models Karpukhin et al. (2020) which separately encode query and passage information, models that see query and passage information jointly at inference time Reimers and Gurevych (2019); Nogueira et al. (2020) are shown to be effective for zero-shot retrieval Rosa et al. (2022). Among those, formulating reranking as sequence generation, such as conducting listwise sorting Ma et al. (2023); Sun et al. (2023b); Pradeep et al. (2023a) or generating rationales Ferraretto et al. (2023), has shown an advantage in application to zero-shot retrieval by leveraging the language model’s auto-regressive generation capabilities. Specifically, a series of studies that use the encoder-decoder architecture of T5 (Sec. 2.2), and applying zero-shot reranking with LLMs (Sec. 2.3), or viewing reranking as autoregressive text generation problem Wang et al. (2024) has been successful.
@@ 2.3 Listwise Reranking with LLMs
2.4 Listwise Reranking with $m$-ary Units
Ai et al. (2019) is in a similar spirit to ours, as they define a basic ranking unit on $m$ elements and extends it to rank the list of $n(\gg m)$ elements. However, while they consider an intractable number of all $m!$ orderings of all $m$-sized subsets and then employ Monte Carlo sampling for approximation, we use $m$-ary tournament sort algorithm which is complete and efficient by construction (Sec. 3.2).
@@ 3.1 Basic Operating Unit ($\boldsymbol{m\to r}$)
The basic operating unit of ListT5 processes a fixed number of $m$ input passages and outputs top-$r$ passages ($m\to r$), as in Fig. 2. We develop the basic unit upon the Fusion-in-Decoder (FiD) architecture Izacard and Grave (2021), which modifies the Encoder-Decoder structure of T5 Raffel et al. (2020) to process listwise inputs efficiently. Specifically, given a query $\mathbf{q}$ and a list of $m$ candidate passages $[\mathbf{p}_{1},...,\mathbf{p}_{m}]$, our basic unit concatenates each passage $\mathbf{p}_{i}$ with its index identifier $i$ and the query $\mathbf{q}$ in the following format, and then feeds them to Encoder to obtain their full sequence token embedding $\mathbf{h}_{i}$:
@@ 3.2 Extension with Tournament Sort ($\boldsymbol{n\to k}$)
Table 1: Comparison of the computational complexity of different ranking methods on obtaining top-$k$ passages from $n$ candidate passages. The base of the log function is $m$. ListT5 achieves $\mathcal{O}(n+k\log{n})$, or $\mathcal{O}(n+n\log{n})$ when $n$ = $k$, which is way better than pairwise models, competitive with pointwise models, and is in fact the best possible (worst-case asymptotic) complexity for sorting algorithms based on comparisons Ren et al. (2018) [^2] https://en.wikipedia.org/wiki/Comparison_sort.
@@ 4.1 Overview
In this section, after explaining the training and evaluation details, we present three main results in the following order: (1) General performance and efficiency comparison of our models with respect to pointwise and listwise ranking models at Sec. 4.4, (2) Analyzing the lost-in-the-middle problem by measuring the robustness to positional bias between ListT5 and listwise ranking models at Sec. 4.5, and (3) Ablation experiments on the design choice at Sec. 6. All results are based on the T5-base model except explicitly mentioning 3B.
@@ 4.2 Training
To train ListT5, we use the official train subset of the MS MARCO Bajaj et al. (2018) passage ranking dataset. [^3] https://microsoft.github.io/msmarco/Datasets The dataset consists of 532,761 distinct queries and 8.8M passages, with binary annotations of relevancy. Since the ordering of relevancy between negative passages is not provided, we use a dual-encoder retrieval model, specifically COCO-DR large Yu et al. (2022), to retrieve Top-1000 passages and label relevance scores of negatives for the training dataset of MS MARCO. Experiments using GTR Ni et al. (2022), the same model used to build the training data for Zhuang et al. (2022), are reported at Appendix. B.2. For each query-positive passage pair, we randomly sample 20 groups of $m-1$ distinct negative passages. We randomly shuffle the positive and negative passages and assign identifiers $\{1,...,m\}$ to them to form each training data.
@@ 4.3 Evaluation
We measure the evaluation performance on the official dev set of MS MARCO (passage ranking dataset), TREC-DL19, and TREC-DL20 Craswell et al. (2021) for in-domain performance, and the BEIR benchmark Thakur et al. (2021) for zero-shot out-domain performance. BEIR contains 18 diverse sets of domains and tasks. We download the dumped index of the Top-100 and Top-1000 retrieved passages by BM25 from the official Pyserini repository. Reranking results using advanced first-stage retrieval models (e.g., COCO-DR) are reported at App. 10. We use the evaluation metric of the Normalized Discounted Cumulative Gain@10 (NDCG@10) Järvelin and Kekäläinen (2002), the official evaluation metric for BEIR Thakur et al. (2021). All reported scores are rounded to the nearest tenth. We also report the Mean Reciprocal Rank@10 (MRR@10) scores at App. Tab. 11. For reproducibility, we include the evaluation details including details to run baseline models and links for dataset download at App. G.
@@
Efficiency in FLOPs. In addition to comparing the computational complexity in Sec. 3.2, we also conduct real-time measurements of floating point operations (FLOPs) using the FlopsProfiler of the DeepSpeed library [^4] https://deepspeed.readthedocs.io/en/latest/flops-profiler.html. We measure the FLOPs needed to rerank Top-10 passages from BM25-Top100 candidate passages for TREC-DL19 (43 queries, maximum input length of 256) on T5-base, run the profiler, and report the relative FLOPs when we set the FLOPs of MonoT5 (pointwise model) as 1555The exact value of FLOPs for MonoT5 in this setup was 229,732,539,806,400.. We also report the performance and FLOPs of the sliding window variants of ListT5, with details at App. 8. The results in Fig. 4 shows that ListT5 ($r$=1) and ($r$=2) are comparable with MonoT5 and RankT5 (pointwise ranking models), and much more efficient than DuoT5 (pairwise ranking models). Also, tournament sort uses output caching to lower down FLOPs than the sliding window variants, with better performance, on the real-time FLOPs calculation (App. 8). Additionally, we can parallize computation of the same level (e.g., leaf node) within query for tournament sort, while the sliding window approach should be computed sequentially. Even better, the ListT5 model is flexible - according to their needs, users can have the choice to select between the trade-off of inference speed ($r$=1) and performance ($r$=2), an option not available for other models.
4.5 Robustness to Positional Bias
Background.
Baseline Models.
Experiment and Data Setup.
4.6 Impact of initial ordering
Initial
ordering
DL19 DL20
TREC-
COVID
TREC-
NEWS
Touche
-2020
Avg.
ListT5-base (tournament sort, r=2)
No shuffle 71.8 68.1 78.3 48.5 33.4 60.0
Shuffle 71.2 68.1 77.2 48.9 32.8 59.6
Perf. drop -0.4
ListT5-base (sliding windows, stride=3, iter=4)
No shuffle 71.8 67.7 77.5 50.0 33.1 60.0
Shuffle 69.5 65.5 77.7 49.2 32.1 58.8
Perf. drop -1.2
RankVicuna-7b (sliding windows)
No shuffle 68.9 66.1 80.5 46.9 33.0 59.1
Shuffle 67.1 64.6 79.2 45.3 30.8 57.4
Perf. drop -1.7
RankGPT-3.5 (sliding windows)
No shuffle 68.4 64.9 72.6 46.5 38.2 58.1
Shuffle 62.5 57.0 66.1 38.3 22.8 49.3
Perf. drop -8.8
Table 5: NDCG@10 results before and after randomly shuffling the initial top-100 ordering of BM25. Evaluation results for RankGPT-3.5 are run 3 times and averaged to compensate for the unstability of the API output. While RankGPT-3.5 suffers heavily after shuffling (-8.8 drop in performance), ListT5 with tournament sort is relatively robust to the initial order shuffling.
All listwise reranking methods, including the tournament sort and the sliding window approach, depends on the initial ordering returned by the first-stage retriever (e.g., BM25). This may leave a room for ordering sensitivity. To investigate this, we experiment on DL19, DL20, and the selected subset of BEIR, to measure the robustness of initial ordering. For each dataset, we shuffle the initial ordering of candidate passages on 3 different seeds (seed 0,1,2) and report the average score. The results from Table 5 show that (1) ordering sensitivity is indeed prevalent in previous listwise reranking models, and that (2) using ListT5 with FiD greatly reduces the sensitivity in general (in line with our findings in Section. 4.5), and (3) When using ListT5, tournament sort is in general more robust to ordering change (-0.4 drop in average performance), compared to sliding windows (-1.2 drop). This means that ListT5 with tournament sort can perform in a robust way even in conditions where the initial ordering is not very trustable, which is mostly the case for zero-shot retrieval tasks.
4.7 Ablation Study
@@ 5 Conclusion
We proposed ListT5, a FiD-based listwise reranking model that jointly considers multiple passages as input and generates an ordered list of passages, in increasing order of relevance. ListT5 outperforms others in BEIR, with a remarkable $+$1.3 gain of average NDCG@10 performance. ListT5 is efficient, leveraging tournament sort to its advantage. Finally, we show that ListT5 overcomes the lost-in-the-middle problem in LLMs, and are more robust regardless of the initial ordering of the passage.
@@ 6 Limitations
We believe that the efficiency of our models can be greatly optimized using simple ideas. For example, if we can implement early stopping at sequential decoding (assuming that the model outputs permutation of 5), we can reduce the decoding step by 80% for ListT5 ($r$=1) and 60% For ListT5 ($r$=2). There could also be other options, such as designing a more efficient way that uses the tournament sort with fewer number of forward passes, efficient way of encoder output caching, and extension to other langauges or other architectures.
@@
7 Acknowledgements
We greatly appreciate the members from the EXAONE lab from LG AI research for providing GPU resources to conduct experiments along with the helpful feedbacks. We would also like to thank the members from the SNU Language & Data Intelligence lab for giving constructive feedbacks to improve the paper. We also thank the anonymous reviewers from ARR December and ARR February for pointing out important discussion points and additional experiments that help support our claim. Lastly, we would like to thank Jinwoo Kim from School of Computing, KAIST for helpful discussions.
@@ B.1 Impact of the number of passages $m$ ListT5 sees at once
MS MARCO 40.7 40.7 - - 40.5 40.7
+ Top-1000 44.7 44.9 - - 44.6 45.0
TREC-DL19 71.2 71.8 - - 70.1 70.5
TREC-DL20 67.3 68.1 - - 66.9 67.2
TREC-COVID 76.7 78.3 78.1 78.3 76.2 77.9
NFCorpus 35.5 35.6 35.3 35.6 36.2 36.6
BioASQ 57.2 56.4 55.9 55.8 55.4 56.4
NQ 52.0 53.1 52.2 52.2 51.5 52.5
HotpotQA 72.1 72.6 71.6 71.6 71.4 71.9
FiQA-2018 39.5 39.6 39.6 39.5 39.0 38.9
Signal-1M (RT) 33.3 33.5 33.3 33.2 31.7 32.0
TREC-NEWS 47.9 48.5 49.2 48.6 47.1 47.8
Robust04 52.0 52.1 51.6 51.7 52.2 53.1
Arguana 49.7 48.9 49.4 49.9 38.6 46.6
Touche-2020 34.2 33.4 33.8 34.2 32.4 32.7
CQADupStack 38.4 38.8 38.5 38.6 38.2 28.8
Quora 86.1 86.4 86.0 86.1 85.5 86.8
DBPedia 43.9 43.7 43.3 43.1 42.7 43.6
SCIDOCS 17.2 17.6 17.4 17.5 17.2 18.0
FEVER 77.8 79.8 78.6 78.4 76.7 79.1
Climate-FEVER 22.8 24.0 23.4 23.3 22.7 23.9
SciFact 74.1 74.1 73.8 74.3 73.4 74.2
Avg(In-domain) 56.0 56.4 - - 55.5 55.9
Avg(BeIR) 50.6 50.9 50.6 50.7 49.3 50.0
@@
This aligns with Sec. 6 from the main paper. In addition to varying the output format, we try to increase the number of passages $m$ the model sees at once, with variants on $r$ as well (See Sec. 3.1 for basic notation). The results from Tab. 7 show that setting $m$ as 10 is not as effective as setting $m$ as 5. We conjecture that there is a trade-off between the (1) benefits from seeing more negatives at once, and (2) the complexity of the ordering task itself. If we increase $m$, the model can get more information, but at the same time, the task becomes very hard, and the boundary of negative order becomes very weak (and not very informative). Therefore, in the case of T5-base models, we find that 5 performed better than 10 passages. Compared to the original FiD for QA Izacard and Grave (2021) where they use large $m$ (=100), ListT5 works best with smaller $m$. We conjecture the reason for this is that sorting task has more complexity than the original task that only use the input as knowledge source, leading to increased effectiveness when $m$ is smaller. Where the use case of the original FiD are to find relevant information among inputs, ListT5 needs to order between different inputs, a task that needs to analyze the relative relevancy between all given inputs. However, we believe the optimal number can be improved (or can be different) for models with different architectures or with larger parametric size.
@@ B.2 Impact of the negative selection model
Table 8: Comparison of ListT5 models trained with GTR and COCO-DR. For ListT5, the reported scores are evaluated using the (r=2) variant. (Sec. B.2)